OpenStudy (moongazer):

Determine the equivalent slope of the line based on the given conditions. 1.)parallel to the line whose equation is x = 2. 2.)perpendicular to the line whose equation is y =3 Please answer clearly and explain the solution clearly and easy to understand. Thanks! (guaranteed medal for a good answer)

OpenStudy (anonymous):

For part b. a line is perpendicular is m2=-1/m1 We have y =3 where m1=0 since y=0*x+3 The line perpendicular is m2=-1/m1=-1/0 which is undefined So y-y1=m(x-x1) y-y1=-1(x-c)/0 0*(y-y1)=-1(x-c) 0=-x+c x=c where c can be any number

OpenStudy (moongazer):

so my answer for 1.) is " any real number" and for 2.) is "m=undefined" am I right?

OpenStudy (anonymous):

For part a. a line is parallel if m2=m1 We have x=2 where m1=1/0 since y=(x-2)/0 The line parallel is m2=m1=1/0 which is undefined So y-y1=m(x-x1) y-y1=1(x-c)/0 0*(y-y1)=-1(x-c) 0=-x+c x=c where c can be any number

OpenStudy (anonymous):

That's another way to do it moongazer :)

OpenStudy (moongazer):

What exactly should I write on my paper?

OpenStudy (anonymous):

pretty much just what I wrote above, do you have any questions about any steps?

OpenStudy (moongazer):

Thanks for answer all my questions. You helped me a lot. :)